OpenStudy (anonymous):

How do you solve 27 to the 1/3 power?

OpenStudy (cwrw238):

power 1/3 = cube root 27 = 3*3*3

OpenStudy (cwrw238):

nope 3^(3 * 1/3) = 3^1 = 3

OpenStudy (cwrw238):

so the solution is 3 u mind a little error as we all do sometimes

OpenStudy (anonymous):

\[\left( 27 \right)^{1/3} = \left( 3^{3} \right)^{1/3} = 3\]

OpenStudy (cwrw238):

rught